OpenStudy (anonymous):

Evaluate: lim (4+h)^2-16 / h h->0

OpenStudy (amistre64):

expand it out if its helpful

OpenStudy (anonymous):

that's what i'm having trouble with

OpenStudy (amistre64):

\[16+h^2 +8h-16 \over h\]

OpenStudy (angela210793):

How abt Hopital's theorem?

OpenStudy (amistre64):

\[\frac{h^2}{h}+\frac{8h}{h}\]

OpenStudy (amistre64):

\(h +8=8\) when h=0

OpenStudy (anonymous):

Where does the 8h come from?

OpenStudy (anonymous):

from opening the (4+h)^2

OpenStudy (amistre64):

(4+h)^2 4+h 4+h ------ 16 +4h +4h +h^2 ------------ 16 +8h +h^2